# Developer Perspectives on Licensing and Copyright Issues Arising from Generative AI for Software Development

### v3

@@ 5.2.3. Fair use and training data
While some developers, as described above, believed that the use of code as training data requires a license, others expressed in open-ended responses that they considered such use to be a fair use, even if it was not always clear whether respondents were referring to this concept in its legal sense or a more general sense. For example, $R_{173}$ said, “I believe that AI training utilizing any published content is a form of fair use and should be encouraged rather than litigated. The usage is entirely transformative rather than derivative.”
Some respondents who believed that using content as training data was fair use compared the practice to the process through which humans produce new works based on the knowledge they have accumulated over time. These respondents believed that it should be assumed that publicly available code is free to use for training, and the burden should be on developers to indicate otherwise. For example, $R_{26}$ said, “Creators need to make efforts to protect data that should not be used in AI training.”
@@
Finding 15: Some developers believed that using code as training data should be considered fair use by default and that the onus should be on developers to protect code that they do not wish to be included in training data.
@@ 5.3. AI-generated output that resembles existing code
Developers also had a range of views on whether it is wrongful for generated code to closely resemble existing code. Some respondents compared what GenAI models do when producing code to what humans have always done (manually) in creating derivative work. $R_{103}$ stated: “Being an open-source software developer, I’m accustomed to seeing my codes copied or ‘cloned’ by other developers without any notice, and I don’t feel anything special from it.” $R_{265}$ noted, “What AI is doing now was being done by developers [throughout] the years through forums, [S]tack [O]verflow, [G]ithub, etc.”
@@
For other respondents, their view on the wrongfulness of generated code that resembled code they had developed depended on the nature of the code being copied, an intuition that tracks U.S. legal doctrine (17 U.S.C. §102, b). $R_{183}$ stated that “small pieces of code are rarely valuable in themselves. For any given function there’s only so many ways to write it. It’s the entire thing that’s valuable.” $R_{246}$ provided a concrete example, saying, “There are only so many ways you can create a fetch request in Java[S]cript.” $R_{573}$ summarized the concept: “Development is typically a process of piecing together various snippets that are customized to a particular use case; there is no expectation of ‘ownership,’ at least in the case of partial code blocks. While the reproduction of an entire application/code base for use in generating revenue is problematic, the sharing of common code is how the development community has progressed [until] now.” And $R_{445}$ offered a useful analogy: “[A]dding a license to a particular piece of code does not make sense. You can add licenses to libraries and exact forms of algorithms. Adding a license to a piece of code is like adding a copyright to individual strokes and colors in a painting.”
For some, the extent of the copying was important, perhaps reflecting U.S. copyright law’s concept of de minimis copying, which does not constitute infringement. $R_{144}$ put it this way: “I feel like if significant pieces of code are copied, only then does it really become an issue for me: I would not be comfortable if someone took multiple files from my code and used them in a proprietary project.” $R_{143}$ said, “If a generative AI happens to have copied large swathes of code, yes [the developers of that code deserve attribution].”
@@ 5.4. On the ownership of AI-generated code
Our discussion to this point has focused on developers’ views on potential liability for copying another’s code, either in training data or in the generated output. Another important legal question is who would own any copyright to code created by GenAI models. In this subsection, we do not attempt to provide a legal analysis of this question; rather, our goal is to understand what developers think the answer to the question should be.
Participants were presented with a list of potential options as to who owns the copyright to code generated through the use of AI and were invited to select all that they believed applied, with the opportunity to provide other options and to further explain their answers. Even though courts have yet to resolve this question, only 57 of 554 respondents familiar with GenAI tools (10%) indicated that they were not sure who owned the output or had no opinion on the matter. The rest of the respondents expressed an opinion and reported high confidence in their answers: 416 (75%) stated they were confident or even very confident.
@@ 5.4.1. No one—works are in the public domain
The most popular choice, selected by 242 participants (44%), was that AI-generated code belongs to no one. $R_{516}$ offered a summary: “Code generated through the use of AI is not copyrightable, because it fails the creative step required by law of being a product of the intellect. Instead it is produced by a non-sentient machine, lacks a creative step, an expression of the soul. It is therefore in the public domain.”
@@
Several respondents invoked the inherently collaborative process of generating code through AI as a reason not to assign copyright ownership to only one individual or entity. $R_{74}$ gave an analogy: “It’d be like saying who gets the copyright of what a child creates. [T]heir teachers? [T]heir parents? [T]heir friends? [T]hey all influence it in an unpredictable way. At least to my knowledge[,] we can’t measure it.” Putting it in more concrete terms, $R_{105}$ stated that the output of GenAI models is “a product of complex interactions between AI algorithms, training data, and user inputs, making traditional copyright assignment impractical. This perspective aligns with the open-source ethos, favoring communal access and innovation over individual ownership in the rapidly evolving field of AI and software development.”
This sentiment regarding the nature of collaborative work was also echoed in responses, particularly from open-source developers, that seemed to suggest that even if the developer could claim copyright in AI output, they should refrain from asserting their rights in favor of the greater good. $R_{233}$ asserted that the focus should be maximizing utility: “I’ve been contributing to open-source code for over a decade, consistently opting for the least restrictive licenses. I firmly believe in the value of sharing knowledge rather than imposing limitations. With the emergence of AI as a lasting tool, my stance is to focus on maximizing its utility for people rather than getting bogged down in minor details.”
Finding 18: The largest group of developers (242, 44%) thought that the output of GenAI belongs in the public domain. Some believe this because it is consistent with the collaborative nature of open-source software development, while others believe that because AI models are not human, the output of those models cannot be subject to copyright law.
@@ 5.4.2. The prompting developer or their employer
The next most popular suggestion was that model output belonged to the developer that prompted the model or their employer, as selected by 201 respondents (36%).
Many participants who reached this conclusion believed that using code-generation tools is similar to existing processes that can result in copyrightable code. Respondents with this perspective, like $R_{173}$, characterized GenAI as “a tool like any other, used to augment the skills and direction of its user,” where the assumption would be that the user owned the rights to whatever was created using the tool. $R_{22}$ provided an analogy: “[I]f a carpenter chooses to use a powerful automation tool like [a] CNC machine to make something, the final product is his (or his employer’s), not the maker of the tool or [whoever] inspired the carpenter.” Ultimately, $R_{118}$ indicated, “If I can copyright code I learned by essentially copying what I learned from Stack Overflow, I should be able to copyright what I get from […] Copilot. […] If I can’t, then the analysis should be similar. The focus on AI as the issue seems misguided. The only difference is scale.”
@@ 5.4.3. The creators whose works make up the training data
Some respondents, by contrast, invoked the collaborative nature of development as a reason to deny copyright ownership claims by developers of the code used in the training data. $R_{157}$ noted, “[T]he neural networks require so much input for them to work that none of the input is really deterministic to the end work. […] So the copyright can not belong to the people that just provided the training data. [If that were true], the copyright of [a] book could belong also to the authors of all the books that the author had read before.” $R_{157}$ continued by providing another analogy from the creative arts: “[We don’t give] copyright to the people that cleaned the theater and people that create the music[al] instruments and so on.”
@@
Finding 20: Some developers believed that fairness requires giving ownership of generated code to those developers who created the training data. However, others invoked the collaborative nature of development to deny copyright claims by those whose code was used in the training data and highlighted the practical problems with mapping training data to generated code to determine ownership.
@@ 5.4.4. The creators of the model
By contrast, other respondents, such as $R_{306}$, indicated that while the ownership question is complicated, model creators were the one group with no real claim: “Again, it is difficult to be sure. What I do believe, however, is that it should never belong to the creators of the model, as they have no say on the output of the model, only its inner workings.” $R_{98}$ expressed a similar sentiment: “The creators of the model are definitely not the copyright owners—LLMs are just text generators and as such can technically generate any text. You can force [a] LLM to ‘write’ literally any sequence of words you want—that [output] can’t be copyrighted [by the model].”
@@ 5.4.5. Other ownership claims
Participants who selected multiple options conveyed that the answer to this question is often case-specific. Previous work that explores how changes to AI supply chains can influence the answers to copyright questions (Lee et al., 2023a) supports this view. $R_{182}$ provided scenarios in which the ownership of generated code might vary: “Code generated by a model that I’ve prompted with a simple prompt (e.g., ‘Please show me a solution to the Fizz-Buzz challenge’) should probably be public domain. However, code generated based on a prompt that includes significant copyright content (e.g., ‘Consider the style and techniques in the 50kB source file below, and then using those styles and relevant techniques, implement a new routine to [do something]. Before you start ask me any clarifying questions you think are important. […]’) could reasonably be protected as if it had been written by a human.” $R_{152}$ wrote, “[It] depends on the jurisdiction, and on how much of the code was generated by the AI, and on how much the output was transformed by the prompter.”
A few respondents chose multiple options because they believed that copyright ownership should be distributed among multiple groups. For example, $R_{191}$ called generated code “a co[-]creation where we would have a percentage of ownership to apply. […] The prompt, however, [would] belong entirely to [the] user.” $R_{527}$ suggested, “I would opt [for] […] a shared benefits model, like [with] royalties in the entertainment industry, where the actual code owner gets a fair share, the model owner, and the end user who was capable of adding the right prompt into the model […].” Such views were uncommon, and no respondent proposed a system for determining an appropriate allocation of ownership rights or the implications thereof.
Finding 22: Some respondents believed that ownership depended on various factors, including the relevant jurisdiction, the amount of code generated, and how much the prompting developer contributed to the result. Others believed that multiple parties share claims to ownership and that a royalty-style system should be put into place.
@@ 5.5. On the ownership and copying of prompts
Much of the current GenAI litigation, as of this writing, relates to the unauthorized use of preexisting work in training data or the copyright status of generated content. We were, however, also interested in understanding developers’ views on the ownership of prompts, which, depending on their complexity, could be protected by copyright law. (As noted above, however, platforms’ Terms of Service might govern users’ rights in this regard.) About half of the 554 respondents familiar with GenAI tools (278, 50%) were indifferent as to whether a prompt they developed could be used by someone else without their permission, with 88 respondents (16%) indicating that they would be pleased if such use occurred and 63 (11%) extremely pleased. An additional 89 respondents (16%) would be displeased by the use of their prompts, and 36 (7%) extremely displeased. Overall, as shown in Figure 8, most developers in our survey stated that they would not care (or would even be pleased) if their prompts were reused by others, even without permission.
@@ 5.6. Perceived risk of potential copyright infringement
Ultimately, in our follow-up interviews, five of seven respondents perceived that the likelihood of legal enforcement resulting from improper usage of GenAI was low. $RI_{39}$ stated in an interview: “If I technically am violating somebody’s copyright for code that is never going to be pushed anywhere public, it’s very implausible that I would be sued for that.” $RI_{31}$ noted that they belonged to “a fairly small organization. We’re not high on the radar. I don’t think that we’re at any legal risk of anything. If anybody’s going to get sued, it’s not going to be us. It’s going to be other people.” $RI_{149}$ expressed a similar view: “[T]he likelihood of detection is so small that it’s no longer a key concern.”
@@ 6. RQ3: Other legal concerns that developers anticipate as the use of generative AI increases
The legal issues other than those related to copyright law that GenAI raises are apparent. RQ3 aimed to discover which of these concerns were top-of-mind among the developers in our study. Developers identified a broad range of issues, including those relating to tort liability, malicious use, and data privacy concerns. Findings 25-29 answer this research question, a summary of which is also available in our replication package (ano, 2024).
@@ 6.1. Liability concerns
Respondents mentioned concerns about possible tort liability (legal responsibility for harmful actions) related to the use of GenAI. Product liability (the claim, for example, that a product is defective or harmful) was mentioned by 14 respondents (3%), and liability for code generated with bugs or defects was mentioned by 13 respondents (2%). In particular, developers were concerned about who would be held accountable for security issues or other bugs introduced by AI-generated code. $R_{0}$ expressed that they “could see in the near future running [into] issues with security breaches caused by an AI code generator making sub-optimal choices. A contrived example[,] but imagine someone uses the AI to build a log-in. The AI suggests that you store passwords in plain text in your database which is later leaked. The developer could argue the model creator (e.g. [O]penAI) was responsible for the breach.”
Other respondents identified potential issues with AI agents, such as consumer-facing chatbots that purport to engage in contractual agreements with users (Belanger, 2024; Malley, 2023) or AI agents that appear to operate autonomously, such that parties may attempt to disclaim responsibility for harmful effects. $R_{149}$ said that “[f]unction calling”—using LLMs “to write and call functions”—was a potential concern (see examples in (AutoGPT, [n. d.]; gptengineer, [n. d.])) and asked, “[W]hat happens if those called (which may have real world effects) cause harm—who is responsible?” Respondents also highlighted the risk posed by using GenAI in safety/mission-critical systems and domains, such as aviation, security, defense, and medical fields.
Finally, 13 respondents (2%) were concerned that code might be generated without sufficient consideration of compliance issues, such as “ignoring GDPR [the EU’s General Data Protection Regulation]” ($R_{190}$). $R_{665}$ was concerned that “[i]f an AI generates specifications for a project, it may not consider existing laws and [so] generate software which does not comply with regulations.” Several respondents raised the related concern of using GenAI to draft legal documents such as software licenses, terms of service, and privacy policies, which could not only lead to an undesirable proliferation of such documents but, if not subject to sufficient human review, also may result in language that does not have the desired legal effect.
Finding 25: Developers raised questions about who would be held accountable for bugs and other defects in GenAI outputs, especially for mission-critical systems such as medical, aerospace, and security. Developers were also concerned about who would be held responsible for actions that AI agents take autonomously.

### v5

@@ 5.2.3. Fair use and training data
While some developers, as described above, believed that the use of code as training data should comply with license requirements, others expressed in open-ended responses that they considered such use to be a fair use, even if it was not always clear whether respondents referred to this concept in its legal sense or in a more general sense. For example, $R_{173}$ said, “I believe that AI training utilizing any published content is a form of fair use and should be encouraged rather than litigated. The usage is entirely transformative rather than derivative.”
Some respondents who believed that using content as training data was fair use compared the practice to the process through which humans produce new works based on the knowledge they have accumulated over time. These respondents believed that it should be assumed that publicly available code is free to use for training and that the burden should be on developers to indicate otherwise. For example, $R_{26}$ said, “Creators need to make efforts to protect data that should not be used in AI training.”
@@
Finding 15: Some developers believed that using code as training data should be considered fair use by default and that the onus should be on developers to protect code from being included in training data.
@@ 5.3. AI-generated output that resembles existing code
Developers also had a range of views on whether it is wrong for generated code to closely resemble existing code. Some respondents compared what GenAI models do when producing code to what humans have always done (manually) in creating derivative work. $R_{103}$ stated: “Being an open-source software developer, I’m accustomed to seeing my codes copied or ‘cloned’ by other developers without any notice, and I don’t feel anything special from it.” $R_{265}$ noted, “What AI is doing now was being done by developers [throughout] the years through forums, [S]tack [O]verflow, [G]ithub, etc.”
@@
For other respondents, their view of the wrongfulness of generated code that resembled code they had developed depended on the nature of the code being copied, an intuition that tracks U.S. legal doctrine (17 U.S.C. §102, b). $R_{183}$ stated that “small pieces of code are rarely valuable in themselves. For any given function, there’s only so many ways to write it. It’s the entire thing that’s valuable.” $R_{246}$ provided a concrete example, saying, “There are only so many ways you can create a fetch request in Java[S]cript.” $R_{573}$ summarized the concept: “Development is typically a process of piecing together various snippets that are customized to a particular use case; there is no expectation of ‘ownership,’ at least in the case of partial code blocks. While the reproduction of an entire application/code base for use in generating revenue is problematic, the sharing of common code is how the development community has progressed [until] now.” And $R_{445}$ offered a useful analogy: “[A]dding a license to a particular piece of code does not make sense. You can add licenses to libraries and exact forms of algorithms. Adding a license to a piece of code is like adding a copyright to individual strokes and colors in a painting.”
For some, the extent of copying was important, perhaps reflecting U.S. copyright law’s concept of de minimis copying, which does not constitute infringement. $R_{144}$ put it this way: “I feel like if significant pieces of code are copied, only then does it really become an issue for me. I would not be comfortable if someone took multiple files from my code and used them in a proprietary project.” $R_{143}$ said, “If a generative AI happens to have copied large swathes of code, yes [the developers of that code deserve attribution].”
@@ 5.4. On the ownership of AI-generated code
Our discussion to this point has focused on developers’ views on potential liability for copying another’s code, either in training data or in the generated output. Another important legal question is who would own any copyright to code created by GenAI models. In this subsection, we do not attempt to provide a legal analysis of this question; instead, our goal is to understand what developers think the answer to the question should be.
Participants were presented with a list of potential options as to who owns the copyright to code generated through the use of AI and were invited to select all that they believed applied, with the opportunity to provide other options and to further explain their answers. Although courts have yet to resolve this question, only 57 of 554 respondents familiar with GenAI tools (10%) indicated that they were not sure who owned the output or had no opinion on the matter. The rest of the respondents expressed an opinion and reported high confidence in their answers: 416 (75%) stated that they were confident or even very confident. Figure 9 shows respondent views on the ownership of model output.
@@ 5.4.1. No one—works are in the public domain
The most popular choice, selected by 242 participants (44%), was that AI-generated code belongs to no one. $R_{516}$ offered a summary: “Code generated through the use of AI is not copyrightable, because it fails the creative step required by law of being a product of the intellect. Instead, it is produced by a non-sentient machine, lacks a creative step, an expression of the soul. It is therefore in the public domain.”
@@
Several respondents invoked the inherently collaborative process of generating code through AI as a reason not to assign copyright ownership to only one individual or entity. $R_{74}$ gave an analogy: “It’d be like saying who gets the copyright of what a child creates. [T]heir teachers? [T]heir parents? [T]heir friends? [T]hey all influence it in an unpredictable way. At least to my knowledge[,] we can’t measure it.” Putting it in more concrete terms, $R_{105}$ stated that the output of GenAI models is “a product of complex interactions between AI algorithms, training data, and user inputs, making traditional copyright assignment impractical. This perspective aligns with the open source ethos, favoring communal access and innovation over individual ownership in the rapidly evolving field of AI and software development.”
This sentiment regarding the nature of collaborative work was also echoed in responses, particularly from open-source developers, that seemed to suggest that even if developers could claim copyright in AI output, they should refrain from asserting their rights in favor of the greater good. $R_{233}$ asserted that the focus should be maximizing utility: “I’ve been contributing to open-source code for over a decade, consistently opting for the least restrictive licenses. I firmly believe in the value of sharing knowledge rather than imposing limitations. With the emergence of AI as a lasting tool, my stance is to focus on maximizing its utility for people rather than getting bogged down in minor details.”
Finding 18: The largest group of developers (242, 44%) thought that the output of GenAI belongs in the public domain. Some align with this view because it is consistent with the collaborative nature of open-source software development, while others believe that because AI models are not human, the output of those models cannot be protected by copyright law.
@@ 5.4.2. The prompting developer or their employer
The next most popular suggestion was that the model output belonged to the developer that prompted the model or their employer, as selected by 201 respondents (36%).
Many participants who reached this conclusion believed that using code-generation tools is similar to existing processes that can result in copyright-compatible code. Respondents with this perspective, like $R_{173}$, characterized GenAI as “a tool like any other, used to augment the skills and direction of its user,” where the assumption would be that the user owned the rights to whatever was created using the tool. $R_{22}$ provided an analogy: “[I]f a carpenter chooses to use a powerful automation tool like [a] CNC machine to make something, the final product is his (or his employer’s), not the maker of the tool or [whoever] inspired the carpenter.” Ultimately, $R_{118}$ indicated, “If I can copyright code I learned by essentially copying what I learned from Stack Overflow, I should be able to copyright what I get from […] Copilot. […] If I can’t, then the analysis should be similar. The focus on AI as the issue seems misguided. The only difference is scale.”
@@ 5.4.3. The creators whose works make up the training data
Some respondents, on the contrary, invoked the collaborative nature of development as a reason to deny copyright ownership claims by developers of the code used in the training data. $R_{157}$ noted, “[T]he neural networks require so much input for them to work that none of the input is really deterministic to the end work. […] So, the copyright cannot belong to the people [who] just provided the training data. [If that were true], the copyright of [a] book could belong also to the authors of all the books that the author had read before.” $R_{157}$ continued by providing another analogy from the creative arts: “[We don’t give] copyright to the people that cleaned the theater and people that create the music[al] instruments and so on.”
@@
Finding 20: Some developers believed that fairness requires giving ownership of generated code to those developers who created the training data. However, others invoked the collaborative nature of development to oppose the idea of copyright claims by those whose code was used in the training data and highlighted the practical problems with mapping the training data to generated code to determine ownership.
@@ 5.4.4. The creators of the model
By contrast, other respondents, such as $R_{306}$, indicated that while the ownership question is complicated, the creators of the models were the only group with no real claim: “Again, it is difficult to be sure. What I do believe, however, is that it should never belong to the creators of the model, as they have no say on the output of the model, only its inner workings.” $R_{98}$ expressed a similar sentiment: “The creators of the model are definitely not the copyright owners—LLMs are just text generators and as such can technically generate any text. You can force [a] LLM to ‘write’ literally any sequence of words you want—that [output] can’t be copyrighted [by the model].”
@@ 5.4.5. Other ownership claims
Participants who selected multiple options conveyed that the answer to this question is often case-specific. Previous work exploring how changes in AI supply chains can influence the answers to copyright questions (Lee et al., 2023a) supports this view. $R_{182}$ provided scenarios in which the ownership of generated code might vary: “Code generated by a model that I’ve prompted with a simple prompt (e.g., ‘Please show me a solution to the Fizz-Buzz challenge’) should probably be public domain. However, code generated based on a prompt that includes significant copyright content (e.g., ‘Consider the style and techniques in the 50kB source file below, and then using those styles and relevant techniques, implement a new routine to [do something]. Before you start ask me any clarifying questions you think are important. […]’) could reasonably be protected as if it had been written by a human.” $R_{152}$ wrote, “[It] depends on the jurisdiction, and on how much of the code was generated by the AI, and on how much the output was transformed by the prompter.”
Some respondents chose multiple options because they believed that copyright ownership should be distributed among multiple groups. For example, $R_{191}$ called generated code “a co[-]creation where we would have a percentage of ownership to apply. […] The prompt, however, [would] belong entirely to [the] user.” $R_{527}$ suggested, “I would opt [for] […] a shared benefits model, like [with] royalties in the entertainment industry, where the actual code owner gets a fair share, the model owner, and the end user who was capable of adding the right prompt into the model […].” Such views were uncommon, and no respondent proposed a system for determining an appropriate allocation of ownership rights or the implications thereof.
Finding 22: Some respondents believed that ownership depended on various factors, including the relevant jurisdiction, the amount of code generated, and how much the prompting developer contributed to the result. Others believed that multiple parties should share ownership claims and that a royalty-style system should be put in place.
@@ 5.5. On the ownership and copying of prompts
Much of the current GenAI litigation, as of this writing, is related to the unauthorized use of preexisting work in training data or the copyright status of generated content. However, we were also interested in understanding developers’ views on the ownership of prompts, which, depending on their complexity, could be protected by copyright law. (As noted above, however, platforms’ Terms of Service might govern users’ rights in this regard.) Approximately half of the 554 respondents familiar with GenAI tools (278, 50%) were indifferent as to whether a prompt they developed could be used by someone else without their permission, with 88 respondents (16%) indicating that they would be pleased if such use occurred and 63 (11%) extremely pleased. An additional 89 respondents (16%) would be displeased by the use of their prompts, and 36 (7%) would be extremely displeased. In general, as shown in Figure 8, most of the developers in our survey stated that they would not care (or even be pleased) if their prompts were reused by others, even without permission.
@@ 5.6. Perceived risk of potential copyright infringement
In our follow-up interviews, five out of seven respondents perceived that, ultimately, the likelihood of legal enforcement resulting from the improper use of GenAI was low. $RI_{39}$ stated in an interview: “If I technically am violating somebody’s copyright for code that is never going to be pushed anywhere public, it’s very implausible that I would be sued for that.” $RI_{31}$ noted that they belonged to “a fairly small organization. We’re not high on the radar. I don’t think that we’re at any legal risk of anything. If anybody’s going to get sued, it’s not going to be us. It’s going to be other people.” $RI_{149}$ expressed a similar view: “[T]he likelihood of detection is so small that it’s no longer a key concern.”
@@ 6. RQ3: Other legal concerns that developers anticipate as the use of generative AI increases
The legal issues other than those related to copyright law that GenAI raises are apparent. RQ3 aimed to discover which of these concerns were top of mind among the developers in our study. Developers identified a wide range of issues, including those relating to tort liability, malicious use, and data privacy concerns. Findings 25-29 answer this research question, a summary of which is also available in our replication package (ano, 2025).
@@ 6.1. Liability concerns
Respondents mentioned concerns about possible tort liability (legal responsibility for harmful actions) related to the use of GenAI. Product liability (the claim, for example, that a product is defective or harmful) was mentioned by 14 respondents (3%), and liability for code generated with bugs or defects was mentioned by 13 respondents (2%). In particular, developers were concerned about who would be held responsible for security issues or other bugs introduced by AI-generated code. $R_{0}$ expressed that they “could see in the near future running [into] issues with security breaches caused by an AI code generator making suboptimal choices. A contrived example[,] but imagine [that] someone uses the AI to build a log-in. The AI suggests that you store passwords in plain text in your database which is later leaked. The developer could argue [that] the model creator (e.g., [O]penAI) was responsible for the breach.”
Other respondents identified potential issues with AI agents, such as consumer-facing chatbots that appear to engage in contractual agreements with users (Belanger, 2024; Malley, 2023) or AI agents that appear to operate autonomously, such that parties may attempt to disclaim responsibility for harmful effects. $R_{149}$ said that “[f]unction calling”—using LLMs “to write and call functions”—was a potential concern (see examples in (AutoGPT, [n. d.]; gptengineer, [n. d.])) and asked, “[W]hat happens if those called (which may have real world effects) cause harm—who is responsible?” Respondents also highlighted the risk posed by using GenAI in safety/mission-critical systems and domains, such as aviation, security, defense, and medical fields.
Finally, 13 respondents (2%) were concerned that code might be generated without sufficient consideration of compliance issues, such as “ignoring GDPR [the EU’s General Data Protection Regulation]” ($R_{190}$). $R_{665}$ was concerned that “[i]f an AI generates specifications for a project, it may not consider existing laws and [so] generate software which does not comply with regulations.” Several respondents raised the related concern of using GenAI to draft legal documents such as software licenses, terms of service, and privacy policies, which could not only lead to an undesirable proliferation of such documents but may also, if not subject to sufficient human review, result in language that does not have the desired legal effect.
Finding 25: Some developers raised questions about who would be held accountable for bugs and other defects in GenAI outputs, especially for mission-critical systems such as medical, aviation, and security. Some developers were also concerned about who would be held responsible for actions that AI agents take autonomously.